Question 16 1 pts The Tiny Tots Toy Company manufactures only sleds. In 2012 Tiny Tots manufactured 10.000 sleds, but sold only 8,000 sleds. In 2012 Tiny Tots change in inventory was 0-2000 sleds. O 1000 sleds. O 2000 sleds O 3000 sleds D | Question 17 pts Assume that in Scandia, planned investment is $80 billion but actual investment is $60 billion. Unplanned inventory investment is O-$20 billion. O -$10 billion. O $70 billion.
Answer #1

Ans.16 - (C) 2000 sleds

Change in inventory = Sleds produced - Sleds sold = 10000-8000=2000

Ans.17- (A) -20

Unplanned inventory investment = Actual investment - Planned investment = 60-80= -20